OpenStudy (anonymous):

I NEED YOUR HELP!!PLEASE :) A line passes through (3, 7) and (6, 9). Which equation best represents the line?

OpenStudy (anonymous):

A line passes through (3, 7) and (6, 9). Which equation best represents the line? y = 3 over 2 x + 5 y = 2 over 3 x + 5 y = 3x + 2 y = 2 over 3 x + 2 here is the choices :)

OpenStudy (princeharryyy):

y = 2 over 3 x + 5

OpenStudy (princeharryyy):

slope = 9-7 / 6-3 = 2/3

OpenStudy (princeharryyy):

y =2/3 x + c put in (3,7) find c

OpenStudy (princeharryyy):

which comes out to be5 @brigamezz

OpenStudy (anonymous):

Ohhh I see that makes more sense :). Do you think you could help me with one more?

OpenStudy (anonymous):

Which equation best represents the line? y = 1 over 3. x − 1 y = 3x − 1 y = −x + 1 over 3. y = 3x + 1

OpenStudy (anonymous):

Which equation best represents the line? A. y = 1 over 3. x − 1 B. y = 3x − 1 C. y = −x + 1 over 3. D. y = 3x + 1

OpenStudy (princeharryyy):

y = 3x − 1

OpenStudy (princeharryyy):

put (-1,-4) it satisfies the equation

OpenStudy (princeharryyy):

@brigamezz